OpenStudy (anonymous):

280 grams of a radioactive substance decays to 100 grams after 500 years. To the nearest year, what is the half-life of the substance?

OpenStudy (anonymous):

you can use \[(\frac{100}{280})^{\frac{t}{500}}=\frac{1}{2}\] solve for \(t\) although you might have been taught differently

OpenStudy (paxpolaris):

\[\Large A=A_0\left( \frac12 \right)^{t/H}\] if H is the half life here... \[\Large 100=280\left( \frac12 \right)^{500/H}\]

OpenStudy (anonymous):

thanks